# Pavlok-Nudge: A Feedback Mechanism for Atomic Behaviour Modification with Snoring Usecase

### v3

@@ Abstract
This paper proposes a feedback mechanism to change behavioural patterns using the Pavlok device. Pavlok utilises beeps, vibration and shocks as a mode of aversion technique to help individuals with behaviour modification. While the device can be useful in certain periodic daily life situations, like alarms and exercise notifications, the device relies on manual operations that limit its usage. To automate behaviour modification, we propose a framework that first detects targeted behaviours through a lightweight deep learning model and subsequently nudges the user through Pavlok. Our proposed solution is implemented and verified in the context of snoring, which captures audio from the environment following a prediction of whether the audio content is a snore or not using a 1D convolutional neural network. Based on the prediction, we use Pavlok to nudge users for preventive measures, such as a change in sleeping posture. We believe that this simple solution can help people to change their atomic habits, which may lead to long-term health benefits. Our proposed real-time, lightweight model (99.8% less parameters over SOTA; 1,278,049 $\rightarrow$ 1337) achieves SOTA performance (test accuracy of 0.99) on a public domain benchmark. The code and model are publicly available at https://github.com/hasan-rakibul/pavlok-nudge-snore.
Index Terms:
@@
Pavlok, behaviour, deep learning, health and wellbeing, snore.
I Introduction
@@
Software-based Behavioural intervention technologies (BITs) have the potential to transform human atomic behaviour [1], which could further lead to long-term health benefits that directly impact human cognition and emotional states [2]. The emerging field of BITs in affective computing mainly includes websites, mobile applications and Internet of Things (IoT) sensor-based designs aiming to make people’s lives better [3, 4]. To this end, we leverage the Pavlok [5], a wearable device aimed to help people ‘break their behavioural pattern’ using aversive conditioning. It accomplishes this by providing nudges in the form of beeps, vibration and/or electric shock to the wearer’s wrist when they engage in targeted behaviours, for example, snoring. The main intention behind this approach is that over time the wearer will become conscious of the atomic behaviour with the unpleasant sensation of the beep/vibration/shock. Even over time, the individual will be less inclined to repeat the behaviour. Moreover, Pavlok can also work in synchronisation with a smartphone application to monitor progress by setting up behaviour change goals. While the native Pavlok is generally useful in certain periodic daily life situations, like alarms, exercise notifications and nail-biting, the device mostly relies on manual operation by the user rather than automatic action detection, which limits its usefulness. Similarly, it lacks the advanced sensory capabilities to detect and take action on relevant events in real time.
@@
Figure 1: Pavlok-Nudge Overview. Pavlok-Nudge is an interactive, nearly real-time software application aiming to aid individuals in atomic behaviour modification. Here, we present the use case for snoring behaviour modification. Our proposed solution consists of the wrist-borne Pavlok and a smartphone (see 1, 2, 3). The microphone on the mobile device captures audio, which is further passed through a deep learning-based framework to detect snoring sounds. Upon snoring detection, the mobile device provides a signal to the Pavlok to nudge the user for posture change to stop snoring (see 2). Scenarios (1), (2) and (3) happen sequentially over time.
@@
Motivation. Snoring is the harsh sound that occurs during sleep when the throat muscles of a person relax, causing narrowness in the airway, which triggers snoring as the relaxed airway vibrates. According to medical research, a natural way to reduce or avert snoring is to sleep on the side [6]. However, it is difficult to maintain the posture involuntarily while a person sleeps. Thus, we believe that our proposed solution, Pavlok-Nudge, can help individuals with behaviour training. Statistically, approximately 75% of people who snore develop obstructive sleep apnea [7, 6]. Sleep apnea leads to the risk of several health issues such as heart disease, diabetes, obesity, arrhythmia (irregular heart rhythm), headache and stroke [8]. Poor quality of sleep reduces memory, thinking capability and proper mental functioning in daily life situations. Prolonged compromises in sleep can lead to frequent irritability, short-temperateness, and even depression [8, 9].
@@
For our use case, the mobile phone’s microphone is used to capture snoring noises. The phone can establish a Bluetooth Low Energy (BLE) connection with the Pavlok device. In this wireless pathway, the phone then sends a signal to the Pavlok device to deliver a ‘nudge’ or ‘shock’ if snoring is detected. Our proposed solution augments the sensory capabilities in the Pavlok. The proposed solution also allows greater customisation and control over Pavlok’s responses related to context. For example, a quiet snore may elicit a vibrate, but a loud snore a shock, and so on. Users can switch between beep, vibration and shock, along with the intensity of the nudge stimulus. Thus, overall, it is user-friendly and adjustable to best suit each individual’s needs. The main contribution of the paper is as follows:
@@
  • Pavlok-Nudge is a software-based behaviour intervention technology to help individuals to break behavioural patterns. It is an interactive, real-time, open-sourced solution that will help the research community to extend it to other human behaviour modification settings.
@@
TABLE I: Where we stand now?: Overview of deep learning algorithms for snoring detection.
@@
Ref. Year Dataset Details Avail. Feature Classifier
[10] 2015 # Sub: 15 $\times$× Raw f-MLP
[11] 2017 # Sub: 68, Lab $\times$× Freq. MLP
[12] 2018 # Sub: 20, Hospital $\times$× MFCC RNN
[13] 2019 # Sub: 10, Bedroom $\times$× Raw 1D-CNN
[14] 2019 # Sam: 1000, Web MFCC 2D-CNN
[15] 2020 # Sub: 15, Hospital $\times$× Mel-spec. CNN-LSTM-DNN
[16] 2021 # Sub: 38, Lab $\times$× CQT-Spec. CNN-LSTM
[17] 2022 # Sam: 1000 Raw NA
[18] 2023 # Sub: 88, Hospital $\times$× Raw 1D/2D-CNN
Our 2024 [17, 14] MFCC 1D-CNN
@@
Figure 2: Pavlok-Nudge Workflow. The Pavlok-Nudge framework senses the environment using the microphone of a mobile or laptop device. To process the audio input, a non-overlapping sliding window-based approach divides the input audio into 1-second chunks, which are then fed to a deep learning-based snore detection model. After collecting 10 such chunks’s output (i.e., 10 seconds of audio), a voting strategy is used to determine if the incoming signal belongs to the snore category. Finally, upon reaching a decision, the system utilises a WiFi signal to connect with the Pavlok device and deliver a user-specific ‘nudge’. For additional details about the workflow, please refer to Section III.
II Related work
@@
Snore Detection and Localisation. Prior literature on snore detection using machine learning methods mainly focuses on detection aspects without any intervention strategy (See Table I for more details). The earliest version of snore recording and detection software was developed by SnoreLab [19]. However, the interface does not provide a snore prevention strategy to act upon snore detection. Non-deep learning-based methods such as logistic regression classifier with auditory image modelling features [20] and AdaBoost classifier with spectral features [21] were utilised for snore detection. Recently, Romero et al. [22] and Emoto et al. [23] use latent space feature from a pre-trained deep learning model to predict if the incoming audio signal is a snore. Similarly, Arsenali et al. [12] use a recurrent neural network with mel-frequency cepstral coefficients (MFCC) features to detect snoring. The most closely related work by Khan [14] also utilises a convolutional neural network (CNN)-based deep learning framework with MFCC feature to detect snoring audio. The proposed solution also includes a vibration device to alert the user to change sleeping posture. However, the deployment situation consists of wired signals, which causes uncomfortable situations for the user. To this end, our solution is more user-friendly and easily deployable.
Pavlok for ‘Breaking Behavioural Pattern’. We leverage the advantage of Pavlok to alert users to change behavioural patterns. The design of the Pavlok device alerts the user periodically or in prior defined time gaps. However, snoring and relevant behaviours require real-time monitoring. For instance, a person who is sleeping cannot pre-define their snoring ahead of time to generate a nudge. To this end, we develop a software solution equipped with the ability to detect and respond to such behavioural events automatically. With the integration of our behaviour-specific deep learning model, the Pavlok is now able to aid individuals in modifying atomic behaviour, particularly which occurs involuntarily.
@@
III Pavlok-Nudge Pipeline
Our proposed ‘Pavlok-Nudge’ solution consists of a deep learning-based snore-detection model, mobile application and Pavlok (See Figure 2).
III-A Deep Learning-Based Snore Detection Model
Input. Given the continuous sensing of audio using the mobile’s microphone, we use a non-overlapping 1-second audio chunk for our analysis. The deep learning-based framework used in our proposed system takes 1-second audio segments to detect the occurrence of snoring incidents. Following the recommendation from previous work [24], we extract 25 MFCC components. For each 1-second audio segment, we compute MFCCs at 85 evenly spaced time points, resulting in a $25\times 8525 × 85$ matrix representation for each audio sample.
Model Architecture. For snore detection, we employ a 1D CNN architecture. After evaluating various configurations of convolutional and fully connected layers, we finalised a model with two convolutional layers, each using $3\times 33 × 3$ kernels, followed by a fully connected layer. Each convolutional layer is followed by a ReLU activation function and max-pooling. The first convolutional layer outputs 8 channels, while the second expands this to 16 channels, allowing for progressively richer feature extraction from the input audio signal. Max-pooling layers with $2\times 22 × 2$ kernels reduce the spatial dimensions by half. After the convolutional layers, the network output is flattened and passed through a fully connected layer. To mitigate overfitting, a dropout layer with a 0.4 rate is applied before the final classification layer. The output is then passed through a Sigmoid activation function to produce a probability for binary classification. This lightweight architecture effectively captures both temporal and frequency features of snore sounds while maintaining a compact footprint of just 1,337 trainable parameters.
Model Training. The model is trained using the Adam optimiser with a learning rate of 0.001 to minimise the binary cross-entropy loss. Training is capped at a maximum of 100 epochs. We tuned various hyperparameters, including the learning rate, dropout rate and convolutional layer sizes, to optimise the architecture. To prevent overfitting and ensure optimal generalisation, we implement an early stopping criterion, halting the training if the validation loss does not improve for three consecutive epochs. The best-performing model, identified by the lowest validation loss during training, is saved to ensure the most effective version is retained for subsequent inferencing.
Dataset. We utilise two snoring datasets. (1) Khan [14] releases a collection of 500 snoring and 500 non-snoring sounds obtained from various online sources. Following the elimination of silent segments from the audio recordings, individual samples lasting one second each are further generated. The snoring sounds are derived from individuals spanning different age groups, including children, adult men and adult women. Among the 500 snoring samples, 72.6% were characterised by the absence of any background noise, while the remaining 27.4% exhibited non-snoring background noise. As for the 500 non-snoring samples, the dataset has encompassed 50 samples for each of the 10 distinct ambient sounds commonly encountered near individuals during sleep, such as a baby crying, clock ticking, toilet flushing, sirens, television noise, car noise, people talking, rain and thunderstorms. Following [14], we split the dataset into three partitions: 700 samples for training, 200 samples for validation and the remaining 100 samples for testing the model.
(2) MaleFemale [17] is another dataset of 1,000 snoring sounds, including 500 from males and 500 from females, is available at Kaggle. As there are no non-snoring sounds in this dataset, we merged this dataset with Khan [14], resulting in a dataset of 2,000 samples. Like Khan [14], we randomly selected 100 samples from the MaleFemale dataset and made a consolidated test set of 200 samples (i.e., 100 from Khan and 100 from the MaleFemale dataset). This merged dataset includes 1,600 training samples (700 from Khan and 900 from the MaleFemale dataset).
Evaluation Protocol. We report the performance of our model in terms of standard classification metrics, including classification accuracy, precision, recall, F1 score and area under the receiver operating characteristic curve (AUROC).
Results. The performance of our snore detection model across different configurations is presented in Table II. Our final model achieved an accuracy of 0.99 on the test set, significantly outperforming the baseline model proposed by Khan [14], which achieved an accuracy of 0.96.
In addition to its superior performance, our model maintains an exceptionally low parameter footprint of only 1,337 trainable parameters. The baseline 2D CNN model requires a much larger and more complex architecture yet achieves a lower accuracy. In addition to better accuracy, our model delivers high precision (0.98), recall (1.00), F1 score (0.99) and AUROC (0.99).
Software Front-End. We developed a software interface that employs JavaScript and Bluetooth technology to facilitate behaviour modification using the Pavlok 3 device [5].
III-B Real-Time Monitoring
The detection cycle in our proposed system comprises the continuous recording of a 10-second audio clip using the microphone of a device such as a smartphone or a computer. This audio clip is then segmented into ten 1-second intervals, which are processed through a pre-trained neural network on the device (smartphone or computer). The neural network is designed to detect the occurrence of snoring incidents in the audio clip. Specifically, if seven out of ten segments are identified as containing snoring incidents, then the app delivers the pre-selected stimulus to the user via the Pavlok device. The detection cycle is designed to continuously monitor snoring behaviour and provide a prompt and effective aversive stimulus to facilitate behavioural modification.
III-C Comparison with Existing BITs
The work more closely related to the proposed approach is the vibration-based feedback mechanism proposed by Khan [14]. They devised a snoring detection model utilising an MFCC feature-based CNN architecture, attaining a commendable accuracy of 0.96. Firstly, our proposed model outperforms this approach in terms of accuracy. Secondly, our approach involves the integration of a custom-made embedded system affixed to the upper arm of individuals, which beeps or vibrates or shocks upon detecting snoring and continues nudging on an adjustably cycle until the person changes their sleeping position, which we could detect using the Pavlok’s accelerometer. Despite its intended functionality, Khan’s device’s bulkiness and wired connections pose discomfort to the user during sleep, which is not a problem with our system.
TABLE II: Performance of our snore detection model and its comparison with the literature.
Data Model O/p chnl #Params Acc. P R F1 AUC
[14]
2D CNN
[14]
$32-$
64-64
1,278,049 0.96
1D CNN

### v5

@@ Abstract
This paper proposes an atomic behaviour intervention strategy using the Pavlok wearable device. Pavlok utilises beeps, vibration and shocks as a mode of aversion technique to help individuals with behaviour modification. While the device can be useful in certain periodic daily life situations, like alarms and exercise notifications, it relies on manual operations that limit its usage. To automate behaviour modification, we propose a framework that first detects targeted behaviours through a lightweight deep learning model and subsequently nudges the user. Our proposed solution is implemented and verified in the context of snoring, which captures audio from the environment following a prediction of whether the audio content is a snore or not using a lightweight 1D convolutional neural network. Based on the prediction, we use Pavlok to nudge users for preventive measures, such as a change in sleeping posture. We believe that this simple solution can help people change their atomic habits, which may lead to long-term health benefits. Our proposed lightweight model (99.8% fewer parameters over SOTA; 790,273$\rightarrow$1,337) achieves SOTA test accuracy of 0.99 on a public benchmark. The code and model are publicly available at https://github.com/hasan-rakibul/pavlok-nudge-snore.
Keywords:
@@
Pavlok behaviour deep learning health and wellbeing aversion therapy snore.
1 Introduction
@@
Software-based behavioural intervention technologies (BITs) have the potential to transform human atomic behaviour [25], which could further lead to long-term health benefits that directly impact human cognition and emotional states [24]. The emerging field of BITs in pervasive computing mainly includes websites, mobile applications and internet of things (IoT) sensor-based designs aiming to make people’s lives better [7, 14]. To this end, we leverage the Pavlok device, a wearable aimed to help people “break bad habits” using aversion therapy [22]. It accomplishes this by providing nudges in the form of beeps, vibration and/or electric shock to the wearer’s wrist when they engage in targeted behaviours, for example, snoring. The main intention behind this approach is that over time, the wearer is expected to become conscious of the atomic behaviour due to the unpleasant sensation of the beep/vibration/shock, which may help the individual to be less inclined to repeat the behaviour. Moreover, Pavlok can also work in synchronisation with a smartphone application to monitor progress by setting up behaviour change goals. While the native Pavlok is generally useful in certain periodic daily life situations, like alarms, exercise notifications and nail-biting, the device mostly relies on manual operation by the user rather than automatic action detection, which limits its usefulness. Similarly, it lacks the advanced sensory capabilities to detect and take action on relevant events in real time.
@@
Figure 1: Pavlok-Nudge Overview. Pavlok-Nudge is an interactive, nearly real-time software application aiming to aid individuals in atomic behaviour modification. Here, we present the use case for snoring behaviour modification. Our proposed solution consists of the wrist-borne Pavlok and a smartphone (see 1, 2, 3). The microphone on the mobile device captures audio, which is further passed through a deep learning-based framework to detect snoring sounds. Upon snoring detection, the mobile device provides a signal to the Pavlok to nudge the user for a posture change to stop snoring (see 2). Scenarios (1), (2) and (3) happen sequentially over time.
@@
Motivation. Snoring is a harsh sound produced during sleep when the throat muscles relax, narrowing the airway and causing it to vibrate as air flows through. According to medical research [29], a natural way to reduce or avert snoring is to sleep on the side. However, maintaining this posture involuntarily could be difficult when a person sleeps. Our proposed system, Pavlok-Nudge, can help individuals with behaviour training. Snoring is strongly correlated with obstructive sleep apnea [16, 1], a condition that causes a person to stop breathing periodically. Sleep apnea causes several health issues such as heart disease, diabetes, obesity, arrhythmia (irregular heart rhythm), headache and stroke [8]. Poor quality of sleep reduces memory, thinking capability and proper mental functioning in daily life situations. Prolonged compromises in sleep can lead to frequent irritability, short temper, and even depression [8].
@@
For our use case, a mobile phone’s microphone is used to capture snoring noises (Fig. 1). The phone can establish a Bluetooth low energy (BLE) connection with the Pavlok device, which receives a signal to deliver a “nudge” or “shock” if snoring is detected. This way, our proposed solution augments the sensory capabilities of the Pavlok. The proposed solution also allows greater customisation and control over Pavlok’s responses related to context. Users can switch between beep, vibration and shock, along with the intensity of the nudge stimulus. Thus, overall, it is user-friendly and adjustable to best suit each individual’s needs. The main contribution of the paper is as follows:
@@
  • Pavlok-Nudge is a software-based behaviour intervention technology to help individuals break behavioural patterns. It is an interactive and open-sourced solution that will help the research community extend it to other human behaviour modification settings.
@@
Table 1: Where do we stand now? Overview of prior studies on snore detection.
@@
Ref. Year Dataset Details Data Avail. Code Avail. Feature Classifier
[17] 2015 # Sub: 15 $\times$ $\times$ Raw f-MLP
[3] 2017 # Sub: 68, Lab $\times$ $\times$ Frequency MLP
[2] 2018 # Sub: 20, Hospital $\times$ $\times$ MFCC RNN
[27] 2019 # Sub: 10, Bedroom $\times$ $\times$ Raw 1D-CNN
[11] 2019 # Sam: 1000, Web $\times$ MFCC 2D-CNN
[10] 2020 # Sub: 15, Hospital $\times$ $\times$ Mel Spectrogram CNN-LSTM-DNN
[28] 2021 # Sub: 38, Lab $\times$ $\times$ CQT Spectrogram CNN-LSTM
[15] 2022 # Sam: 1000 Mel Spectrogram CNN
[13] 2023 # Sub: 88, Home $\times$ $\times$ Raw; Visibility Graph 1D-2D-CNN
[5] 2025 [11] $\times$ MFCC Multi-branch CNN
[12] 2025 [11] $\times$ Mel Spectrogram CNN-GRU
[19] 2025 [11] $\times$ Time-frequency VGG19
[20] 2025 [11] $\times$ Time-frequency VGG16-SVM
Our 2025 [11, 15] MFCC 1D-CNN
@@
Figure 2: Pavlok-Nudge Workflow. The Pavlok-Nudge framework senses the environment using the microphone of a mobile or laptop device. To process the audio input, a non-overlapping sliding window-based approach divides the input audio into 1-second chunks, which are then fed to a deep learning-based snore detection model. After collecting predictions on a certain number of chunks (e.g., 10 seconds), a voting strategy (e.g., 70%) is used to determine if the incoming signal belongs to the snore category. Finally, upon reaching a decision, the system utilises Bluetooth to connect with the Pavlok device and deliver a user-specific “nudge”. Refer to Sect. 3 for additional details or model architecture, such as kernel size, dropout rate and activation functions.
2 Related Work
@@
Table 1 presents an overview of prior work on snore detection using various deep learning methods. One of the earliest versions of snore recording and detection software was developed by SnoreLab [26]. However, the interface does not provide a snore prevention strategy to act upon snore detection. Non-deep learning-based methods, such as a logistic regression classifier with auditory image modelling features [18] and an AdaBoost classifier with spectral features [4], were utilised for snore detection. Recently, Romero et al. [23] and Emoto et al. [6] use latent space features from a pre-trained deep learning model to predict if the incoming audio signal is a snore. Similarly, Arsenali et al. [2] use a recurrent neural network with mel-frequency cepstral coefficients (MFCC) features to detect snoring. The most closely related work by Khan [11] also utilises a convolutional neural network (CNN)-based deep learning framework with MFCC features to detect snoring. The proposed solution also includes a vibration device to alert the user to change sleeping posture. However, their system relies on wired connections that may cause user discomfort, and its deep learning model is relatively large and computationally intensive. Our proposed solution is more user-friendly and lightweight for easier deployment.
3 Pavlok-Nudge Pipeline
Our proposed “Pavlok-Nudge” solution consists of a deep learning-based snore-detection model, mobile application and Pavlok (Fig. 2).
3.1 Deep Learning-Based Snore Detection Model
3.1.1 Problem Formulation.
Let the continuous audio signal captured by the mobile’s microphone be denoted as $x(t)$, where $t$ represents time. The audio is segmented into non-overlapping 1-second chunks, $x_{i}$, such that:
$x=\{x_{1},x_{2},\dots,x_{N}\},\quad\text{where }i=1,2,\dots,N.$ (1)
Each 1-second audio segment $x_{i}\in\mathbb{R}$ is processed to extract MFCCs. As suggested in prior work [9], we compute 25 MFCC components at 85 evenly spaced time points, resulting in a feature matrix $X_{i}\in\mathbb{R}^{25\times 85}$ for each audio segment $x_{i}$.
Thus, the input becomes a sequence of $X=\{X_{1},X_{2},\dots,X_{N}\}$, where $X_{i}$ serves as the representation of the corresponding audio segment $x_{i}$. Our goal is to develop a classifier $\mathcal{F}$ that maps each input audio segment $X_{i}\in\mathbb{R}^{25\times 85}$ to a binary output $y_{i}\in\{0,1\}$, where $y_{i}=1$ indicates the presence of snoring and $y_{i}=0$ otherwise.
3.1.2 Model Architecture.
Our proposed 1D CNN architecture captures local temporal patterns in the MFCC time series with a relatively small parameter footprint, which is crucial for deployment with limited computational resources. After evaluating various configurations of convolutional and fully connected layers, we finalised a model with only two convolutional layers, each using $3\times 3$ kernels, followed by a fully connected layer. Each convolutional layer is followed by a ReLU activation function and max-pooling. The first convolutional layer outputs 8 channels, while the second expands this to 16 channels, allowing for progressively richer feature extraction from the input audio signal. Max-pooling layers with $2\times 2$ kernels reduce the spatial dimensions by half. After the convolutional layers, the network output is flattened and passed through a fully connected layer. To mitigate overfitting, a dropout layer with a 0.4 rate is applied before the final classification layer. The output is then passed through a Sigmoid activation function to produce a probability for binary classification. This lightweight architecture effectively captures both temporal and frequency features of snore sounds while maintaining a compact footprint of just 1,337 trainable parameters.
3.1.3 Evaluation Protocol.
We report the performance of our model in terms of standard classification metrics, including classification accuracy, precision, recall, F1 score and area under the receiver operating characteristic curve (AUROC).
3.2 Real-Time Monitoring
The detection cycle in our proposed system comprises continuous audio recording using the microphone of a device such as a smartphone or a computer. For our experiments, we use a 10-second recording window to balance timely detection with sufficient context. This audio clip is then segmented into 1-second intervals, which are processed through the pre-trained snore detection model $\mathcal{F}$. If a specified proportion of the audio segments (e.g., 70%) is identified as containing snoring incidents, a pre-selected stimulus – such as beeps, vibration or an electric shock – is administered to the user via the Pavlok device, in accordance with their preferences. The detection cycle is designed to continuously monitor snoring behaviour and provide a prompt and effective aversive stimulus to facilitate behavioural modification.
4 Experiments and Results
4.1 Dataset
We utilise two snoring datasets. (1) Khan [11] released a collection of 500 snoring and 500 non-snoring sounds obtained from various online sources. After removing the silent segments from the audio recordings, individual samples lasting one second each were separated. The snoring sounds were derived from individuals spanning different age groups, including children, adult men and adult women. Among the 500 snoring samples, 72.6% were characterised by the absence of any background noise, while the remaining 27.4% exhibited non-snoring background noise. As for the 500 non-snoring samples, the dataset consists of 50 samples for each of the 10 distinct ambient sounds commonly encountered near individuals during sleep, such as a baby crying, clock ticking, toilet flushing, sirens, television noise, car noise, people talking, rain and thunderstorms. Following [11], we split the dataset into three partitions: 700 samples for training, 200 samples for validation and the remaining 100 samples for testing the model.
(2) MaleFemale dataset, collected from Kaggle [15], consists of 1,000 snoring sounds: 500 from males and 500 from females. As there are no non-snoring sounds in this dataset, we merged this dataset with the Khan [11] dataset, resulting in a dataset of 2,000 samples. Like Khan [11], we randomly selected 100 samples from the MaleFemale dataset and made a consolidated test set of 200 samples (i.e., 100 from Khan and 100 from the MaleFemale dataset). This merged dataset includes 1,600 training samples (700 from Khan and 900 from the MaleFemale dataset).
4.2 Implementation Details
The experiments were conducted on a system running SUSE Linux Enterprise Server 15 SP4 with Python 3.12.3 as the programming language. We utilised PyTorch as the core framework, along with torchaudio for audio processing, torchmetrics for evaluation metrics, and pytorch-lightning for efficient model training. Additionally, ffmpeg was used to support audio data processing. We developed the software interface using JavaScript and Bluetooth technology to establish a connection with the Pavlok 3 device [22].
The CNN models are trained using the Adam optimiser with a learning rate of 0.001 to minimise the binary cross-entropy loss. Training is capped at a maximum of 100 epochs. We tuned various hyperparameters, including the learning rate, dropout rate and convolutional layer sizes, to optimise the architecture. To prevent overfitting and ensure optimal generalisation, we implement an early stopping criterion, halting the training if the validation loss does not improve for three consecutive epochs. The best-performing model, identified by the lowest validation loss during training, is saved to ensure the most effective version is retained for subsequent inference.
4.3 Main Result
The performance of our snore detection model across different configurations is presented in Table 2. Our final model achieved an accuracy of 0.99, along with high precision (0.98), recall (1.00), F1 score (0.99) and AUROC (0.99), on the test set. While two recent works [5, 12], published in 2025, also reported competitive performances, their model size is much larger (parameter count of 39.9 and 0.79 million, respectively) with a more complex architecture. In addition to superior performance, our model maintains an exceptionally low parameter footprint of only 1,337 trainable parameters.
Table 2: Performance of our proposed model and its comparison with the literature.
Dataset Model O/p Ch $\downarrow$ Params $\downarrow$ Acc $\uparrow$ P $\uparrow$ R $\uparrow$ F1 $\uparrow$ AUROC $\uparrow$
Khan [11] 2D CNN [11] 32-32-64-64 1.3 M 0.96
Multi-branch CNN [5] 384- $\cdots$ -2144 39.9 M 0.99 1.00 0.99
CNN-GRU [12] 0.79 M 0.99 0.99 0.99 0.99 0.99
VGG19 [19] 64- $\cdots$ -512 144 M 0.97 0.94 0.99 0.96 1.00
VGG16-SVM [20] 64- $\cdots$ -512 138 M 0.96 0.96 0.96 0.96
2D CNNa [11] 32-32-64-64 1.3 M 0.95 0.91 1.00 0.95 0.95
1D CNN (Ours) 8-16 1,337 0.99 0.98 1.00 0.99 0.99
Khan [11] + MaleFemale [15] 1D CNN (Ours) 8-16 1,337 0.96 0.97 0.98 0.97 0.94
  • a
    Our implementation of [11]’s model, evaluated on identical test samples as ours.
  • “–” means unreported.
  • Parameters in italics refer to our calculation based on the architectures reported in the corresponding papers.
  • O/p Ch – number of channels in the output of convolutional layers
  • Acc – classification accuracy; Params – number of trainable parameters; P – precision; R – recall
The pioneering work of Khan [11] proposed an MFCC feature-based CNN architecture and reported a classification accuracy of 0.96. The Khan [11] dataset does not have any fixed, separated test samples. Therefore, to ensure a fairer comparison by evaluating on identical test samples, we reimplemented Khan [11]’s proposed model [^1] Our implementation of [11] achieved a classification accuracy of 0.95, compared to the 0.96 accuracy reported by them. This slight difference is likely due to variations in data preprocessing methods.. As shown in Table 2, our proposed 1D CNN model outperforms [11]’s 2D CNN model in terms of accuracy, precision, F1 score and AUROC. Both Khan [11] and our model provide a recall of 1.00, indicating that it is easier to identify all snore instances in this dataset, with no false negatives.
In terms of the aversion technique, Khan [11] proposed a custom-made embedded system affixed to the upper arm, which uses a vibration-based feedback mechanism. In contrast, our approach leverages the commercially available Pavlok device, which offers a range of feedback options – including beeps, vibrations or shocks – in adjustable nudging cycles. While Khan’s system fulfils its intended functionality, its bulkiness and wired connections may cause discomfort for users during sleep – an issue mitigated by the compact and wireless design of our system.
Fig. 3 depicts the receiver operating characteristic (ROC) curves for the test sets in two configurations: one trained on the Khan [11] dataset and the other trained on the combined Khan [11] & MaleFemale [15] dataset. Although the model trained on the Khan dataset achieves an AUROC of 0.99, the model trained on the combined dataset achieves a slightly lower AUROC of 0.94. This difference suggests that the inclusion of additional data from the MaleFemale dataset may introduce complexities or variations that slightly affect the model’s classification ability.
@@ 4.3 Main Result
4.4 Ablation Study
Table 3 presents the performance evaluation of different 1D convolutional layer configurations, varying in the number of output channels, when tested on the Khan [11] dataset. While three out of four configurations achieve a recall of 1.00, there are slight variations in accuracy, precision, F1 score and AUROC. The model with two 1D convolutional layers having output channels of 8 and 16 achieves the best performance. The smallest model, with two 1D convolutional layers (output channels of 4 and 8), has the fewest parameters (577) and achieves an accuracy of 0.96 and AUROC of 0.96. On the other hand, the largest model (1,873 parameters), with 16 output channels, yields an accuracy of 0.97, precision of 0.94, F1 score of 0.97 and AUROC of 0.97, slightly lower than the 8-16 configuration despite having more parameters. Since the 8-16 configuration strikes the best balance between model size and performance, we select this as the optimal configuration for our proposed system.
Table 3: Performance comparison of 1D convolutional layer configurations on the Khan [11] dataset.
O/p Channels $\downarrow$ Params $\downarrow$ Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ AUROC $\uparrow$